Find a counterexample to the statement “every positive integer greater than 7 can be written as the sum of the squares of three (not necessarily unique) integers.”


1
Expert's answer
2021-11-09T17:09:08-0500

Legendre's three-square theorem states that a natural number can be represented as the sum of three squares of integers "{\displaystyle n=x^{2}+y^{2}+z^{2}}" if and only if "n" is not of the form "{\displaystyle n=4^{a}(8b+7)}"  for nonnegative integers "a" and "b."

The first numbers that cannot be expressed as the sum of three squares (i.e. numbers that can be expressed as "{\displaystyle n=4^{a}(8b+7)})" are


"{\displaystyle n=4^{0}(8(0)+7)}=7"

"{\displaystyle n=4^{0}(8(1)+7)}=15>7"

"{\displaystyle n=4^{0}(8(2)+7)}=23"


"{\displaystyle n=4^{1}(8(0)+7)}=28"

"{\displaystyle n=4^{0}(8(2)+7)}=23"

"{\displaystyle n=4^{0}(8(3)+7)}=31"

"{\displaystyle n=4^{0}(8(4)+7)}=39"

"{\displaystyle n=4^{0}(8(5)+7)}=47"

"{\displaystyle n=4^{0}(8(6)+7)}=55"

"{\displaystyle n=4^{1}(8(0)+7)}=60"

"..."